OpenStudy (anonymous):

Answers for Midpoint And Endpoint A(-3,5) And B(9,8)

OpenStudy (anonymous):

do you want the other endpoint

OpenStudy (anonymous):

if these are endpoints: (x1+x2)/2=x (-3+9)/2=6/2=3 (y1+y2)/2=y (5+8)/2=13/2=6.5 midpoint is (3,6.5)

OpenStudy (anonymous):

if one is a midpoint and one is an endpoint: (x+9)/2=-3 (y+8)/2=5 x=-15 y=2 the other endpoint is (-15,2)